# for Academic Information Seeking

### v1

@@ 5.2 Query Generation
Following semantic augmentation, we replace the special symbols in the templates (represented by uppercase “X”s) with information about a specific academic entity in an academic data soruce. A placeholder in a question template must be replaced by the attributes of specific academic entities, such as person_id, name, interest, organization, bio, educational_experience, etc., as illustrated in Figure 2. Taking the instance of the alma mater query again, The name of this scholar will be used to fill in the $Q_{T}$ to become a real question “What is scholar name’s alma mater” and the educational experience of this scholar serves as the golden answer to this query.
Table 1: The semantics of symbols in $Q_{T}$.
@@
5.3 Code Generation
In the final step, we employ ChatGPT again to generate a piece of code that corresponds to a given combination $C$ with missing input information. This code is intended to get the answer of the template query $Q_{T}$ shared the same combination $C$ with it. Still with information from a specific academic entity, the $i_{1}$ serves as the input of the generated code. The code is expected to be executed to get the same result as $o_{J}$, which can be seen as a golden information of this code. If the result matches the golden information, this code is valid. Any invalid code, not producing matching results, will be excluded from the final dataset.
6 LLMs Alignment
In order to ultimately leverage LLM to accomplish the task of academic information seeking, SoAy needs to align LLM to be capable of calling academic APIs correctly and efficiently through generating solutions and code. This can be achieved by constructing training data and aligning the LLMs according to it, whether through a tuning or tuning-free method.
6.1 SoAyBench–{(Query, Solution, Code)}
Table 2: Question statistics in SoAyBench set.
Question Type One-hop Two-hop Three-hop Total
@@ 6.1 SoAyBench–{(Query, Solution, Code)}
After going through the aforementioned four stages, a substantial amount of high-quality (Query, Solution, Code) data has been automatically collected. We organize this data into a comprehensive benchmark known as SoAyBench. Currently, SoAyBench consists of 3,960 triplets, out of which one-fifth has undergone expert inspection and adjustment to form a test set. This test set is used to evaluate the ability of the LLMs to search academic information using the AMiner API. The remaining portion of the data is utilized to support the alignment of the LLMs. Table 2 shows the question statistics of SoAyBench.
Table 3: Overall Performance comparison of the proposed SoAy and the baseline methods ToolLLaMA and GTP-DFSDT, as tested on SoAyBench. The terms 3.5, 3.5-16K, and 4 are abbreviations for different versions of GPT. 7B and 13B denotes the parameter scale of the LLM. SoAyBench is derived from 44 question templates, we calculate each metric within a group of instantiated questions of these templates, and then report the mean and variance across groups with different instantiated entities.
@@
6.2 Model Alignment
@@
We provide alignment strategy for both super large closed-source LLMs and tunable open-source LLMs. (1) SoAyGPT is our proposed solution for models like GPT Eloundou et al. (2023), which cannot have their parameters modified but possess strong instruction following capabilities. We select a diverse set of triplets from SoAyBench data that represent different usages of solutions as examples to construct prompts. Given a query, SoAyGPT generates the solution and code in sequence through in-context learning, followed by executing the generated code to obtain the result. We utilize different versions of GPT models as backbones for this purpose. (2) SoAyLLaMA, on the other hand, is our alignment strategy for the LLaMA Touvron et al. (2023) model, which is an open-source model that we can fine-tune within an acceptable training cost. By using the query of the triplets from SoAyBench as input and integrating the solution and code as output, SoAyLLaMA is trained to generate the solution and code in a single step based on the given query in a zero-shot setting.
7 Experiments
In this section, we conduct experiments to validate the effectiveness, efficiency, and practicality of SoAy. We focus on the following Research Questions: (RQ1) In which part of the academic information seeking do LLMs struggle the most? (RQ2) In terms of effectiveness on academic information seeking, does SoAy outperform existing methods? (RQ3) In terms of efficiency, does SoAy finish the task within an execution time that users can accept? (RQ4) What kind of base model is more suitable for constructing SoAy?
We conduct experiments on SoAyBench to compare SoAy with other baseline methods, focusing on these research questions. Additionally, we introduced a detailed evaluation method called SoAyEval, which could also be used in the evaluation of other tool-using scenarios. We also perform online evaluations specifically targeting real-world academic information seeking use cases. There’s an independent chapter in the appendix showing more implementation details and interesting findings.
7.1 Experimental Settings
Baselines. Many existing tool-utilizing works fail to function effectively due to API inter-dependency and intricate execution logic. We have chosen GPT-DFSDT, employing the same three versions of the GPT as backbone as SoAyGPT and ToolLLaMA-7b Qin et al. (2023b) as baselines in API-using experiment on SoAyBench, as the former represents the state-of-the-art model on a general domain tool-utilizing benchmark, ToolBench Qin et al. (2023b), and the latter is an open-source model that has been fine-tuned on numerous tool-utilizing data samples, achieving results comparable to GPT-DFSDT. For every LLM, we adopt a temperature setting of 0 to ensure reproducibility.
It is worth noting that the responses from the baselines, ToolLLaMa, and GPT-DFSDT, are in natural language format, whereas the answers in SoAyBench are the responses returned by SoAPI. In order to ensure a fair comparison, we have designed a set of prompts to employ ChatGPT to extract precise answers from responses in a natural language format.
SoAyEval. Given a question that yields an API solution along with an invoked answer, we have outlined five types of evaluation metrics. These metrics take into account the accuracy of the solution, program, and answer together: (EM): Both the retrieved solution and answer Exactly Match the ground truth. (DS): The answer is correct, but a Different Solution is retrieved compared to the ground truth. (WS): The answer is wrong due to a Wrong Solution. (WP): The solution is correct but the answer is wrong, due to a Wrong Program generated for the solution, which can be executed but yields the wrong answer. (EE): Execution Error, which may by caused by the generation of a non-executable program or network errors during the APIs request. Based on the aforementioned metrics, we define the accuracy metric as follows:
[…]
@@ 7.1 Experimental Settings
We name this evaluation method SoAyEval. SoAyEval meticulously analyzes the entire process of large models using a set of APIs to answer corresponding questions. It is a universal evaluation method that could be adopted by other API-using scenarios. Notably, we assign $w_{1}=1,w_{2}=2,w_{3}=3$ in our experiment to highlight the multi-hop question-answering ability of the methods under test.
7.2 Overall Performance (RQ1, 2)
Table 3 presents the experimental results of SoAyGPT, SoAyLLaMA, and the other two representative baseline methods on SoAyBench. SoAyLLaMA achieves the highest overall score, with Code-13B Model of 91.24%, followed by the other SoAy method, SoAyGPT, which gets 86.57% with GPT-4 as the backbone. These results prove the effectiveness of SoAy. With the help of the prepared solution library, SoAy describes the complex inter-dependency and diverse execution logic to the LLMs. This significantly reduces the pressure of LLMs when determining the optimal execution path from the coupling APIs.
Despite undergoing specialized tool usage training, ToolLLaMA does not score highly. This is primarily because ToolLLaMA’s training objective is to master the use of general APIs, where the dependencies involved are significantly less complex than those in complicated academic APIs. Furthermore, many questions in SoAyBench require a solution that incorporates branch or loop logic. Facing with the need to execute these complicated execution logics, GPT-DFSDT’s decision-tree-based inference encounters numerous repetitive sub-nodes, making the solution exceptionally challenging.
7.3 Error Analysis (RQ1)
The detailed analysis of model performances from SoAyEval reveals distinct characteristics and potential improvement areas for each model across different question types. The SoAyLLaMA and SoAyGPT models, particularly their latest versions, demonstrate superior capabilities in managing complex question types with minimal errors. ToolLLaMA and early versions of GPT-DFSDT show greater susceptibility to execution and solution pathway errors, indicating potential areas for algorithmic or training enhancements. Overall, advancements in model versions (notably from GPT-DFSDT 3.5 to 4 and SoAyGPT 3.5 to 4) show tangible improvements in reducing specific error types and enhancing overall accuracy, highlighting the benefits of continuous model development and refinement.
7.4 Efficiency (RQ3)
We evaluate the efficiency of the proposed method by reporting the average response time of various methods evaluated on SoAyBench, as shown in Table 4. We apply both SoAyLLaMA and ToolLLaMA in a network request service format to keep justice with GPT-based methods which should only be accessed through APIs calls. Both ToolLLaMa and GPT-DFSDT rely on multiple executions to complete complex API operations, thus consuming a considerable amount of time. Implementation through SoAy can effectively reduce the inference time of tool calls. Even with GPT-4, the average execution time of SoAy is nearly one-third of the execution time of GPT-DFSDT with GPT-4. Moreover, with a smaller parameter scale, the lightweight SoAyLLaMA models show even less execution time, which we consider already meet the requirement for being applied to an academic searching service.
Table 4: Average response time of different methods (second).
@@ 7.4 Efficiency (RQ3)
7.5 Backbones Comparison (RQ4)
It’s worth considering which backbone model to choose to construct an academic searching system whether through SoAyGPT or SoAyLLaMA. Based on the results of SoAyBench, if we focus on the effectiveness of the results, the choice of SoAyGPT definitely falls on the latest version (GPT-4 when the paper is drafted). As for the choice of SoAyLLaMA, whose implementation is described in Appendix B, we can find a trend where the scores increase as the number of parameters grows. Simultaneously, given that Soay implements the planning and formatting processes defined in the problem section through code generation, CodeLlama, which excels in code generation, outperforms ChatLlama. It should be mentioned that the effectiveness shown on the results of SoAyBench should not be the sole criterion for judgment when considering practical applications.
8 Related Works
APIs, short for Application Programming Interfaces, are the software interfaces that provide services to other software components [^1] https://en.wikipedia.org/wiki/API#cite_note-1. It facilitates the exchange of data between various software applications, thereby enabling interoperability between them Ofoeda et al. (2019). For the application of LLMs using the API-using methodology, the current solutions and technical approaches can be divided into two major groups: A straightforward idea is to broaden the tool/API-using coverage for LLMs. Such efforts not only include basic web services, such as search engines Nakano et al. (2021); Liu et al. (2023); Zhang et al. (2023) and calculators Schick et al. (2023), but also include the intelligent multi-modal functions, such as vision understanding Wu et al. (2023) and image generation Lu et al. (2023).
Others are devoted to the tool’s dependencies, programming logic and planning ability King et al. (2023); Shen et al. (2023); Paranjape et al. (2023). Except for the efforts that regard LLMs as database managers Qin et al. (2023a), there are also plenty of methods that modeling the tool’s graphical relations Zhang (2023), building intermediate instructions Burnett and McIntyre (1995) and accomplish complex tasks Song et al. (2023).
Besides, applying LLMs to directly access the data from the Database through techniques like text-to-SQL Qin et al. (2022) proves ineffective. The limitation arises from the unavailability of access to the original data sources of related systems.
9 Conclusion

### v2

@@ 4.2. Query Generation
Template-Query Instantiation. For each template query $Q_{T}$, we substitute the placeholders denoted by uppercase “X”s with specific attributes such as person_id, name, interest, organization, bio, and educational_experience of academic entities from an academic data source to create the instantiated query $Q$. For instance, using the alma mater query as an example again, the placeholder in $Q_{T}$ is replaced with the actual name of a scholar denoted by $i_{1}^{\prime}$, resulting in a specific question like “What is scholar name’s alma mater?”. The actual educational background of this scholar denoted by $o_{j}^{\prime}$, then serves as the ground truth answer to $Q$.
Table 1. The semantics of symbols in $Q_{T}$.
@@
4.3. Code Generation
So far, a solution $S$ is tailored to a specific intent by associating head API input $i_{1}$ and tail API output $o_{J}$ to form a combination $C=(i_{1},S,o_{J})$, which is linked to three template queries $\{Q_{T}\}$. Each $Q_{T}$ is instantiated with real academic entities, resulting in a set of specific queries $\{Q\}$, each with its own input $i_{1}^{\prime}$ and corresponding ground truth $o_{J}^{\prime}$.
Building on this data, we then utilize ChatGPT to generate a piece of code corresponding to each combination $C$. This code is designed to retrieve answers for various queries that share the same combination. Using $i_{1}^{\prime}$ from a query $Q$ as the input for the code, we execute the code and then compare its output to the ground truth $o_{J}^{\prime}$. If the output aligns with the ground truth, the code is considered valid. Any code that fails to produce matching results is excluded from the final dataset.
5. LLMs Alignment
To align a general-purpose LLM with the mode of generating corresponding solutions and code based on given problems, we need to prepare relative training data (Section 5.1) and choose the aligning method (Section 5.2).
5.1. SoAyBench–{(Query, Solution, Code)}
Table 2. Query statistics of the Created Dataset.
Query Type One-hop Two-hop Three-hop Total
@@ 5.1. SoAyBench–{(Query, Solution, Code)}
After going through the aforementioned four stages, a substantial amount of high-quality (Query, Solution, Code) data has been automatically created. Table 2 shows the query statistics of the dataset. Currently, SoAyBench consists of 17 solutions with 44 corresponding combinations. Each combination’s template query is augmented 3 times and then, through 30 times of query instantiation, results in 90 real queries. This process yields a total of $44\times 3\times 30=3,960$ triplets. One-fifth of these triplets have undergone expert inspection and adjustment to form a test set, which is published as SoAyBench. This test set is supported by a publicly available academic API service, cloned from the online AMiner service. It is used to evaluate the ability of LLMs to search academic information using the AMiner API. The remaining data is used as the training data to support the alignment of LLMs.
Table 3. Overall Performance comparison of the proposed SoAy and the baseline methods ToolLLaMA and GTP-DFSDT, as tested on SoAyBench. The terms 3.5, 3.5-16K, and 4 are abbreviations for different versions of GPT. 7B and 13B denotes the parameter scale of the LLM. SoAyBench is derived from 44 question templates, we calculate each metric within a group of instantiated questions of these templates, and then report the mean and variance across groups with different instantiated entities.
@@
5.2. Model Alignment
We provide alignment strategy for both super large closed-source LLMs and tunable open-source LLMs.
Closed-source LLMs Alignment. SoAyGPT is our proposed solution tailored for models like GPT, which, while unable to have their parameters modified, exhibit strong instruction-following capabilities. We engage these closed-source LLMs through in-context learning, leveraging training data from our created dataset to generate prompts. We primarily create three types of prompts: solution, code, and answer generation prompts.
The solution generation prompt is constructed by integrating all solutions explained by their corresponding template queries from the training data. For each solution in the solution library $L_{A}$, we create a specific code generation prompt by incorporating various instantiated queries and code snippets corresponding to the solution from the training data. Since LLMs may generate solutions not existing in $L_{A}$, we also create a general code generation prompt for any new solution by including instantiated queries and code snippets of all solutions in $L_{A}$. The answer generation prompt includes examples of queries, code execution results, and corresponding natural language answers to curate human-friendly responses.
Upon receiving a query, SoAyGPT first generates a solution using the solution generation prompt. If the solution can be found in $L_{S}$, the specific code generation prompt for it is used to generate the code; otherwise, the general code generation prompt is used. The generated code is then executed to produce the result. Finally, the answer generation prompt is used to compose the query and the execution result into a natural language answer. All these prompts are illustrated in Figure 7 in the Appendix.
@@ 5.2. Model Alignment
Open-course LLMs Alignment. SoAyLLaMA, on the other hand, is our alignment strategy for LLaMA (Touvron et al., 2023), which is an open-source model that we can fine-tune within an acceptable training cost. By using the instantiated query of the triplets from the training data as input and integrating the solution and code as output, SoAyLLaMA is trained to generate the solution and code in a single inference step based on the given query in a zero-shot setting.
6. Experiments
In this section, we conduct experiments to validate the effectiveness, efficiency, and practicality of SoAy. We focus on the following Research Questions: (RQ1) Is the SoAy method more effective than the baseline method in the task of utilizing the AMiner API to assist academic information retrieval? (RQ2) Is the SoAy method more efficient than the baseline method? (RQ3) Can the performance of the SoAy method in practical applications satisfy users?
We first conduct a detailed comparison of the effectiveness of SoAy and previous methods on the SoAyBench test set using our carefully designed metric, SoAyEval. Then, we deploy an application by SoAy and validate its effectiveness on collected user queries.
6.1. Experiments on SoAyBench
6.1.1. Experimental Settings.
We propose a detailed and interpretable evaluation method based on SoAyBench and conduct comparative experiments with representative baselines.
SoAyEval. Enabling large models to use tools for academic information retrieval is typically a multi-step task, with different steps often requiring the model to demonstrate different capabilities. We need detailed evaluation metrics that can help us compare the strengths and weaknesses of different methods and models in specific steps and capabilities. We propose SoAyEval, a detailed evaluation method designed for SoAyBench. Given a (Query, Solution, Code) triplet from SoAyBench, SoAyEval first executes the Code in AMiner Environment to get the ground truth answer, then compares the tested result with these ground truth solutions, code, and answer. (EM): Both the solution and answer Exactly Match the ground truth. (DS): The answer is correct, but a Different Solution is generated compared to the ground truth. (WS): The answer is wrong due to a Wrong Solution. (WC): The solution is correct but the answer is wrong, due to a Wrong Code generated for the solution, which can be executed but still leads to a wrong answer. (EE): Execution Error, which may be caused by the generation of a non-executable code or network error during the request. Based on the aforementioned metrics, we define the accuracy metric as follows:
$\text{ACC}=\text{EM}+\text{DS},$
@@ 6.1. Experiments on SoAyBench
Notably, we assign $w_{1}=1,w_{2}=2,w_{3}=3$ in our experiment to highlight the multi-hop question-answering ability of the methods under test.
Baselines. We select GPT-DFSDT, utilizing the same three GPT versions as SoAyGPT, and ToolLLaMA-7b (Qin et al., 2023b) as baselines. GPT-DFSDT represents the current state-of-the-art method on multiple general domain tool-utilizing benchmarks(Qin et al., 2023b; Guo et al., 2024). ToolLLaMA-7b is an open-source model fine-tuned on a large corpus of tool-utilizing data, achieving comparable results to GPT-DFSDT.
Since there are no benchmark datasets or methods for evaluating the task of academic information retrieval using APIs, we select the two methods mentioned above, which respectively represent the most effective and efficient methods in general domain tool-using tasks. We believe these methods can provide relatively strong comparisons for validating the effectiveness of SoAy in the domain of academic information seeking.
It is worth noting that results from ToolLLaMA and GPT-DFSDT are in natural language format. To ensure a fair comparison using SoAyEval, we instruct ChatGPT to extract precise answers from their responses. Also, ToolLLaMA and GPT-DFSDT don’t have a module to generate a solution or plan before inferring, so we extract the solution from their path of API calls. For all LLMs, a temperature setting of 0 is used to ensure reproducibility.
Table 4. Average response time of different methods (second).
@@
6.1.2. Overall Performance (RQ1).
Table 3 presents the experimental results of SoAyGPT, SoAyLLaMA, and the other two representative baseline methods on SoAyBench. SoAyLLaMA achieves the highest overall score, with Code-13B Model of 92.74%, followed by the other SoAy method, SoAyGPT, which gets 86.57% with GPT-4 as the backbone. These results prove the effectiveness of SoAy. With the help of the prepared solution library, SoAy describes the complex inter-dependency and diverse execution logic of the LLMs. This significantly reduces the pressure of LLMs when determining the optimal execution path from the coupling APIs.
Despite undergoing specialized tool usage training, ToolLLaMA does not score highly. This is primarily because ToolLLaMA’s training objective is to master the use of general APIs, where the dependencies involved are significantly less complex than those in complicated academic APIs. Furthermore, many questions in SoAyBench require a solution that incorporates branch or loop logic. Facing the need to execute these complicated execution logic, GPT-DFSDT’s decision-tree-based inference encounters numerous repetitive sub-nodes, making the solution exceptionally challenging.
6.1.3. Efficiency (RQ2).
We evaluate the efficiency of the proposed method by reporting the average response time of various methods evaluated on SoAyBench, as shown in Table 4. We apply both SoAyLLaMA and ToolLLaMA in a network request service format to keep justice with GPT-based methods which should only be accessed through API calls. Both ToolLLaMA and GPT-DFSDT rely on multiple executions to complete complex API operations, thus consuming a considerable amount of time. Implementation through SoAy can effectively reduce the inference time of tool calls. Even with GPT-4, the average execution time of SoAy is nearly one-third of the execution time of GPT-DFSDT with GPT-4. Moreover, with a smaller parameter scale, the lightweight SoAyLLaMA models show even less execution time, which we consider already meets the requirement for being applied to a real-world service.
Figure 4. Distribution of questions used for online human evaluation. Questions are categorized by 1, 2, and 3-hop types, further divided into exact and ambiguous categories. The ambiguous category is specific to the online evaluation setting.
[…]
Distribution of questions used for online human evaluation. Questions are categorized by 1, 2, and 3-hop types, further divided into exact and ambiguous categories. The ambiguous category is specific to the online evaluation setting.
6.1.4. Variances.
Notably, the data fluctuations shown in Table 3 do not reflect the variations that would result from repeating the same experiment on identical samples. Given that the $768$ test samples in SoAyBench originate from $44$ template questions, and each template is filled with different entities to create unique question instances, we compute each metric within a group of instantiated questions from the $44$ question templates. We then calculate the mean and variance across groups of different instantiated questions. Our observations indicate that SoAyGPT-4 and SoAyLLaMA deliver a more consistently stable performance across various error types, which leverages GPT-4’s improved generalization ability and the high-quality training data provided by SoAyBench.
6.1.5. Intriguing phenomena
Table 3 reveals several other intriguing phenomena: First, while SoAy excels with GPT-4, its advantages are less marked on GPT-3.5-turbo and GPT-3.5-16k (hereafter collectively referred to as GPT-3.5 backbones), trailing behind the GPT-DFSDT method with the same GPT-3.5 backbones. Notably, the DS and WS metrics for SoAy with GPT-3.5 backbones are quite high (the lower, the better). This can be attributed to GPT-4’s superior instruction-following proficiency, which amplifies the impact of SoAy’s complex prompts on the model adept at instruction following. Second, SoAy methods exhibit superior performance on the two-hop questions than one and three-hop questions, particularly with GPT-3.5 backbones. This is likely because the solution parsing prompts sampled from the SoAyBench are biased towards two-hop questions (refer to Table 2 for more information), causing the GPT-3.5 backbones to default to two-hop solutions within the SoAy framework. On the contrary, GPT-4’s robust instruction-following ability enables it to perform consistently across all question types. Third, GPT-DFSDT exhibits a stronger understanding of one-hop questions when using GPT-3.5 backbones. This is because the general idea of GPT-DFSDT involves breaking down complex questions into simpler one-hop questions, determining the appropriate APIs and their parameters step by step, which enhances its comprehension of one-hop questions.
6.2. Deployment Evaluation
Experiments on SoAyBench validate the effectiveness and efficiency of SoAy. However, in practical applications, the questions that users are truly concerned with may differ from those in the dataset. To test the user experience of SoAy in practical applications, we conduct the following deployment experiment.
6.2.1. Settings.
We deploy SoAy as a web application, then gather 52 representative real user queries from the logs, and invite 15 annotators to conduct human evaluation.
Queries Detailes. The distribution of the 52 gathered queries is illustrated in Figure 4. We classify them not only into 1, 2, and 3-hop types as SoAyBench does but also into either exact or ambiguous questions. Ambiguous questions, such as requests for “comparisons between two scholars”, lack standardized answers and thus yield diverse responses.
Figure 5. Voting results for authentic online questions.
[…]
Voting results for authentic online questions.
Baselines. We compare the responses from SoAy, GPT-4, and human expert annotated answers on these 52 questions. Here GPT-4 indicates GPT-4-1106-preview and SoAy indicates SoAyGPT with GPT-4 as the backbone. We do not include GPT-DFSDT, ToolLLaMA, or other tool-using works as comparisons because, based on the results from Section 6.1.3, their efficiency clearly cannot support the requirements of practical deployment.
Metrics. We invite 15 annotators to choose their favorite answer from three for each query. For a balanced and fair assessment of the experimental results, we develop an annotation platform and create a detailed annotation manual. This platform allows annotators to shuffle and anonymize the answers from the three methods before each voting session. This task did not require the annotators to have expert-level knowledge, so we recruit annotators who are interested in academic tasks through public channels. As such, the annotators’ level of diligence and cognitive abilities might vary significantly, and we aim to filter out annotators who did not annotate seriously, or those whose opinions deviate significantly from the mainstream, through consistency analysis. We examine the Spearman correlation coefficients among the voting results of the 15 human annotators, as shown in Figure 6. We sum up all the coefficients of each annotator to obtain a trustworthiness score for this annotator. We then select the top 10 annotators based on this score and use their votes for further analysis. For each question, the method selected by the most participants (out of ten) receives one vote. The final vote totals for each method are calculated by summing the votes across all 52 questions.
Figure 6. Spearman correlation among 15 annotators
6.2.2. Results (RQ3).
The summarized results are presented in Figure 5. These results show a preference for the SoAy method’s responses in real-world scenarios across various question types, outperforming both direct GPT-4 and human expert responses. Notably, SoAy’s responses are particularly favored for questions with exact answers. However, SoAy’s advantage is less pronounced for ambiguous questions when compared to baseline responses. This observation aligns with SoAy’s design and the assessment objectives of SoAyBench, both of which prioritize precise answering through the integration of AMiner API Library resources. Nevertheless, there’s a noticeable gap between the nature of real-world inquiries and the templated questions in SoAyBench, posing a challenge for SoAy in addressing certain scenarios. This finding points towards the potential for integrating API calls with model-intrinsic functions for a more robust response system.
7. Related Works
In the era of Large Language Models (LLMs) (Zhao et al., 2023), the task of academic information retrieval can be accomplished by teaching LLMs to use the APIs of established academic systems (dblp team, 2005; Giles et al., 1998; Noruzi, 2005; Tang et al., 2008; Sinha et al., 2015). Teaching LLMs to handle these APIs is not easy as LLMs’ abilities are limited by the extent of knowledge acquired during pre-training (Radford et al., 2019). For the LLMs API-using methodology (Qin et al., 2023a), the current technical approaches can be divided into two major directions:
A straightforward idea is to broaden the tool or API-using coverage for LLMs. Such efforts have tried to teach LLMs to handle more and more colorful tools/APIs, from the most basic web services (Nakano et al., 2021; Liu et al., 2023b; Zhang et al., 2023; Lazaridou et al., 2022a; Komeili et al., 2021; Lazaridou et al., 2022b) to calculators (Schick et al., 2024; Das et al., 2024), KB accessing tools (Thoppilan et al., 2022; Patil et al., 2023; Hao et al., 2024; Zhuang et al., 2024), code interpreter (Lu et al., 2023; Gao et al., 2023; Pan et al., 2023; Wang et al., 2023), multi-modal functions (Wu et al., 2023; Sur´ıs et al., 2023; Yang et al., 2023a, b; Gao et al., 2024; Zhao et al., 2024b) and other scenarios (Zhao et al., 2024a; Yao et al., 2022; Inaba et al., 2023).
Some other works have focused on the need to connect multiple tools, requiring LLMs to analyze and plan these relationships (King et al., 2023; Shen et al., 2023; Paranjape et al., 2023). Examples include Chameleon (Lu et al., 2023), which firstly retrieves relevant APIs based on the given query and then separately executes them to get the responses, ToolChain (Zhuang et al., 2023), which uses a decision tree for planning, RestGPT (Song et al., 2023), which combines coarse and fine-grained planning, and ControlLLM (Liu et al., 2023a), which reasons based on a graph structure. These works often use prompt-based methods with limited manual examples, restricting LLMs’ reasoning potential, and can be inefficient due to the high number of reasoning steps needed. Training-based methods like ToolLLM (Qin et al., 2023b), ToolLink (Qian et al., 2023), TPTU-v2 (Kong et al., 2023), and OpenAGI (Ge et al., 2023) can teach LLMs to fully understand some APIs as there’s no limitation in the scale of training examples. However, these works focus more on APIs that can be individually used like weather query APIs and calculator APIs, but ignore the coupling between the APIs. Our method, SoAy, uses both tuning-free and training-based approaches to solve the coupling problem in academic APIs to seek academic information.
8. Conclusion